OpenStudy (anonymous):

abs(x-1) < 2(x-1) , solve inequality please!!

OpenStudy (anonymous):

\[\left| x-1 \right| \le 2(x-1)\]

OpenStudy (anonymous):

if \(x>1\) then \(|x-1|=x-1\) solve \[x-1\leq 2x-2\]

OpenStudy (anonymous):

you get \[1\leq x\] you know know it is true if \(x\geq 1\)

OpenStudy (anonymous):

then if \(x<1\) you have \(|x-1|=1-x\) solve \\[1-x\geq 2(x-1)\]